Was meant to be SBI  recheck on colonies to give clearance from EFB.   Due to SBI being off sick it was cancelled at the last minute,  but as we had delayed going through the colonies on Sunday we had to check them over.
Weather :-  HOT  and Sunny,  Little wind,  24 degrees.
Hive#1    not gone through as poss Virgin Queen
Hive#2    not gone through as poss virgin Queen
Hive#4    Inspected by Richard and Liz.   Brood Chamber:   National Deep.  QNS, no Qc’s. Brood and new eggs, room/stores ok.   4 supers,  3 ready to take off.  –  lots of play cups but none charged.  Temper 7
Hive#5   Inspected By Richard and Liz.    Brood chamber: Deep and Half.   QNS,   3 hatched Qc’s.  1 deep frame of brood/eggs . lot of drones – first time gone through for a month.  Temper 7
Hive#6  Not gone through as poss virgin Queen.
Hive#7  Not gone through
Hive# 7A    Inspected by  Richard and Liz,   Brood chamber:  4 Shallows ( Peter Street’s Method)                QNS,  Several play cups, not charged,  5 f bas in top super, room and stores ok. temper 8
Hive#8     Inspected by Richard and Liz.    Brood chamber: Deep.   QNS,  no Qc’s.  2 f BAS,  stores ok but room very limited,  Q laying right up to the frames .   2 supers, both near full, needs another next visit.  Temper  9.
Hive# 15 F,   Inspected by Richard and Liz,  Brood chamber: Deep.    QS,   1 QC,  Knocked down. 6 f BAS,  stores /room ok.  1 super( nearly full) –  Needs another Super next visit. Temper 9
Hive#15 H,    Inspected by Richard and Liz,  Brood chamber: Deep.   QNS,  3 Charged Qc’s knocked down. 6 F BAS with new eggs.  lack of room to lay so made up another Deep and moved another 3 ft towards Hive 3 slot.   Temper: 9.
Nuc’s  16B and 16C  not gone through.